OpenStudy (anonymous):

what is the square root of 8/50

OpenStudy (welshfella):

just grab your calc and find it

OpenStudy (welshfella):

or do you want to simplify it in radical form?

OpenStudy (anonymous):

0.05656854249

OpenStudy (welshfella):

sqrt8 = sqrt4 * sqrt2 - 2 sqrt2 sqrt50 = sqrt2 * sqrt25 = 5 sqrt2 so sqrt 8/50 = 2 sqrt2 / 5 sqrt2 = 2/5